OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is (a) parallel to the given line and (b) perpendicular to the given line. (0,2), y= -4x +6

OpenStudy (mertsj):

y-2=-4(x-0) y-2=1/4(x-0)